OpenStudy (anonymous):

1/ the square root of ab

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{ab} }\]

OpenStudy (anthonyym):

Do you want to rationlize it? Multiply numerator and denominator by \[\sqrt{ab}\] (since it it's essentially multiplying by 1, it doesn't change the value)

OpenStudy (anonymous):

rationalize the denominator

OpenStudy (anonymous):

multiply top and bottom by \(\sqrt{ab}\) then the radical will be out of the denominator (and move to the numerator)

OpenStudy (anthonyym):

Ok so, \[\frac{ 1 }{ \sqrt{ab} }*\frac{ \sqrt{ab} }{ \sqrt{ab} }\]

OpenStudy (anthonyym):

When multiplying fractions, you multiply straight across - the numerator of the first fraction times the numerator of the second fraction, denominator times denominator. You see that \[\sqrt{ab}*\sqrt{ab}=ab\] and \[1*\sqrt{ab}=\sqrt{ab}\]

OpenStudy (anonymous):

i put the square root of ab and i got it wrong

OpenStudy (anthonyym):

Ok let's rewind. Do you know how to multiply fractions? For example \[\frac{ 2 }{ 3 }*\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

if the denominators are not the same then you would have to find the lcd

OpenStudy (anthonyym):

No, that's in ADDING/SUBTRACTING fractions. When you multiply fractions, it's much easier. You just multiply the numerators and denominators straight across.

OpenStudy (anonymous):

so would it be 10/9

OpenStudy (anthonyym):

Yay correct

OpenStudy (anthonyym):

So the point of "rationilizing" the denominator is to make it so it doesn't have a square root. You do that by multiplying by another square root of that same number. See above. sqrt(ab) + sqrt(ab) = ab

OpenStudy (anthonyym):

my bad it should be sqrt(ab) * sqrt(ab) there

OpenStudy (anonymous):

i put ab and thats also wrong

OpenStudy (anthonyym):

And whatever you do to the denominator of the fraction, you also have to do to the numerator of the fraction. So if you multiply the denimantor by sqrt(ab), you also have to multiply the numerator by sqrt(ab)

OpenStudy (anthonyym):

The reason is that \[\frac{ \sqrt{ab} }{ \sqrt{ab} }=\frac{ 1 }{ 1 }=1\], and if you multiply anything by 1 it's still the same value.

OpenStudy (anonymous):

1 is also wrong

OpenStudy (anthonyym):

Please read everything carefully as a whole so you understand it all. Maybe scroll back up.

OpenStudy (anthonyym):

So you should be doing \[\frac{ 1 }{ \sqrt{ab} }*\frac{ \sqrt{ab} }{ \sqrt{ab} }\]